OpenStudy (anonymous):

Find the exponential function f(x) = Ca^x whose graph goes through points (0,3) and (2,12).

OpenStudy (xapproachesinfinity):

hey there

OpenStudy (misty1212):

HI!!

OpenStudy (xapproachesinfinity):

if the graph goes through that point then we can say that f(0)=3

OpenStudy (misty1212):

oh @xapproachesinfinity got this right?

OpenStudy (xapproachesinfinity):

i guess @misty1212 lol

OpenStudy (xapproachesinfinity):

so urena do we have an agreement about what i just wrote ?

OpenStudy (anonymous):

yeah, just say alex not my last name xp

OpenStudy (xapproachesinfinity):

ok i assumed it wasn't your real name sorry :) so we can say f(0)=3 and f(2)=12 yes

OpenStudy (anonymous):

aha its cool, and yass

OpenStudy (xapproachesinfinity):

ok good now we see f(0) another by using that weird exponential function they give us \[f(0)=Ca^0=C\] similarly \[f(2)=Ca^{12}\] yes?

OpenStudy (anonymous):

yeah i see how you got that

OpenStudy (xapproachesinfinity):

alright! now since we have f(0)=3 and f(0)=C then C=3 yes?

OpenStudy (anonymous):

yass

OpenStudy (xapproachesinfinity):

now we found C we know that f(2)=12, and f(2)=Ca^12=3a^2 then 3a^2=12 yes

OpenStudy (anonymous):

i thought it would be 3a^12 = 2

OpenStudy (xapproachesinfinity):

oh i did and error before after i said similarly \[f(2)=Ca^2 \] a^2 not a^12 did a typo there sorry

OpenStudy (anonymous):

ok i get it

OpenStudy (xapproachesinfinity):

yes my mistake, i was evaluating 2 but put 12 unwillingly

OpenStudy (xapproachesinfinity):

ok now \[3a^2=12\] is equation we can solve for a

OpenStudy (xapproachesinfinity):

any ideas how to solve it?

OpenStudy (anonymous):

divide by 3

OpenStudy (xapproachesinfinity):

ok good start a^2=4 what is next

OpenStudy (anonymous):

square both sides so that a cancels and you have 4 and -4

OpenStudy (xapproachesinfinity):

hmm you mean square root both sides

OpenStudy (anonymous):

yeah my bad xp

OpenStudy (xapproachesinfinity):

\[a^2=4 \Longrightarrow a=\pm \sqrt{4}\]

OpenStudy (anonymous):

2 and -2

OpenStudy (xapproachesinfinity):

so a=2 or a=-2 yes

OpenStudy (xapproachesinfinity):

good! now we have to decide which is valid

OpenStudy (anonymous):

how?

OpenStudy (xapproachesinfinity):

what did you learn about exponential function don't they tell all the time a>1 yes?

OpenStudy (anonymous):

right, so a=2

OpenStudy (anonymous):

thank you so much

OpenStudy (xapproachesinfinity):

well a can be less than 1 too but not negative since that can cause problems

OpenStudy (anonymous):

ok note-to-self

OpenStudy (xapproachesinfinity):

so what they actually say is \[0<a<1~~or ~~a>1\]

OpenStudy (xapproachesinfinity):

you are welcome~

OpenStudy (xapproachesinfinity):

thanks for the testimony :)